Problem 10.11 4 Question 7 of 8 D Check My Work 10-2: Basic Definitions 10-5: The Cost of Retained Earnings, rs WACC and Percentage of Debt Financing Hook Industries capital structure consists solely of debt and common equity. It can issue debt at rd 1190, and its common stock currently pays a $3.50 dividend per share (Do $3.50). The stocks price is currently $21.75, its I dividend is expected to grow at a constant rate of 7% per year, its tax rate is 35%, and its WACC is 15.05%, what percentage of the companys capital structure consists of debt? Round your answer to two decimal places.
Answer #1
Cost of debt 7.15% =11%*(1-35%)
Cost of equity 24.22% =3.5*(1+7%)/21.75+7%

WACC = Wd×Rd×(1-t)+We×Ke

W is weights of respective portfolios

R is return on respective portfolios

Wd+We = 1

15.05% = Wd×7.15%+(1-Wd)×24.22%

15.05% = Wd×7.15%+24.22%-24.22%×Wd

17.07%×Wd = 9.17%

Weight of debt, Wd = 0.5372

Hence, percentage of debt is 53.72%